OpenStudy (lena772):

Calculus: Find the derivative by the limit process for f(x)=3x+2

OpenStudy (owlcoffee):

The definition of derivative is: \[\lim_{x \rightarrow a} \frac{ f(x)-f(a) }{ x-a }\]

OpenStudy (lena772):

lim 3x+2-f(a) x-->a x-a

OpenStudy (owlcoffee):

\[\lim_{x \rightarrow a} \frac{ (3x+2)-(3a+2) }{ x-a }\]

OpenStudy (lena772):

Is that it?

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x \rightarrow a} \frac{ (3x+2)-(3a+2) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3x+2-3a-2 }{ x-a }\)

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x \rightarrow a} \frac{ (3x+2)-(3a+2) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3x+2-3a-2 }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3x-3a }{ x-a }\)

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x \rightarrow a} \frac{ (3x+2)-(3a+2) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3x+2-3a-2 }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3(x-a) }{ x-a }\)

OpenStudy (anonymous):

\[f'(x) = \lim_{h \rightarrow 0} \frac{ f(x+h)-f(x) }{ h }\] \[f'(x) = \lim_{h \rightarrow 0} \frac{ (3(x+h)+2)-(3x+2) }{ h }\] \[f'(x) = \lim_{h \rightarrow 0} \frac{ 3x+3h-3x }{ h }\] \[f'(x) = \lim_{h \rightarrow 0} \frac{ 3h }{ h }\] \[f'(x) = \lim_{h \rightarrow 0} 3\] \[f'(x) = 3\]

ganeshie8 (ganeshie8):

\(\large \lim \limits_{x \rightarrow a} \frac{ (3x+2)-(3a+2) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3x+2-3a-2 }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3(x-a) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a}~~~ 3\)

ganeshie8 (ganeshie8):

now that bottom cannot equal 0, you can substitute x = a \(\large \lim \limits_{x \rightarrow a} \frac{ (3x+2)-(3a+2) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3x+2-3a-2 }{ x-a }\) \(\large \lim \limits_{x \rightarrow a} \frac{ 3(x-a) }{ x-a }\) \(\large \lim \limits_{x \rightarrow a}~~~ 3\) \(\large 3\)

ganeshie8 (ganeshie8):

Also, you canfirm that 3 is correct answer by seeing that given funciton is simply a straight line wid slope of 3 :- f(x)=3x+2 ^